a rectangle has length measuring twice it width. if it width is 7cm what is its perimeter
Answer by macston(5194)

W=width=7cm; L=length=2W; P=perimeter=2(L+W)
L=2W=2(7cm)=14cm
P=2(14cm+7cm)
P=2(21cm)
P=42cm ANSWER: The perimeter is 42 centimeters.